OpenStudy (anonymous):

PLEASE HELP! Solve for x using square root property and simplify: (4x+1)^2=8

OpenStudy (solomonzelman):

square both sides, and subtract 1 from both sides.

OpenStudy (solomonzelman):

I mean square root both sides...

OpenStudy (anonymous):

Eh...was going to draw it out, but my hand eye coordination is terrible. Basically what Solo says. Root both sides to cancel the square. You'll end up with 4x+1=the root of 8. Then subtract 1. After, divide by 4.

OpenStudy (anonymous):

That'll give you x.

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf (4x+1)^{2}=8 }\) \(\Huge\color{blue}{ \bf \sqrt{(4x+1)^{2}}=± \sqrt{8} }\) \(\Huge\color{blue}{ \bf \sqrt[2]{(4x+1)^{2}}=± \sqrt{8} }\) \(\Huge\color{blue}{ \bf \sqrt[\cancel{ 2} ]{(4x+1)^{\cancel{ 2}}}=± \sqrt{8} }\) \(\Huge\color{blue}{ \bf 4x+1=± \sqrt{8} ~~~.... }\)